Hi, I know how to do a proportion but I'm not totally sure how to set up this question and what it is asking, here's the problem: You have a photograph that is 4 inches wide and 5 1/3 inches tall. You would like to reproduce it on a sheet of paper that measures 8 1/2inches by 11inches, leaving at least 1/2 inch margin on all four sides. What scale should you use if you keep the sheet of paper in the normal upright orientation. Assume that the reproduction will be as large as possible
Answer by stanbon(75887)

You have a photograph that is 4 inches wide and 5 1/3 inches tall. You would like to reproduce it on a sheet of paper that measures 8 1/2inches by 11inches, leaving at least 1/2 inch margin on all four sides. What scale should you use if you keep the sheet of paper in the normal upright orientation. Assume that the reproduction will be as large as possible
----
The usable dimensions on the finised sheet are 7.5 by 10
----
Solve:: 4x = 7.5
x = 15/8 or 1.875
-----
(5 1/3)x = 10
(16/3)x = 10
x = 30/16
x = 15/8 = 1.875
-----------------------
Ans: Scale = 1.875 or 15/8
-------------
Cheers,
Stan H.
------------